OpenStudy (anonymous):

What mass of sulfur combines with 8 grams copper to form copper (I) sulfide Cu2S?

OpenStudy (abmon98):

1)Construct a balanced chemical equation 2Cu+S-->Cu2S 2)Number of moles=Mass(g)/Molar Mass(g/mol) 3) according to the balanced chemical equation, prepare a mole ratio of the reactant and products. Use again Number of moles=Mass(g)/Molar Mass(g/mol) but rearrange it Number of moles*Molar Mass(g/mol)=Mass(g) Look for coppers and sulphur atomic weight in your periodic table.

OpenStudy (anonymous):

Thanks!

OpenStudy (abmon98):

your most welcome :)